OpenStudy (anonymous):

WILL MEDAL AND FAN! Perform the following operations and write the answers in rational exponent form. √7 + √3 + √98 - √18 and 3√5 - 3√11 + 2√121 -3√90

OpenStudy (anonymous):

@SerenityJustice74

OpenStudy (anonymous):

@leonardo0430

OpenStudy (serenityjustice74):

Sorry, I was messging my friend. Okay what do you think? Where can you start? Think about PEMDAS

OpenStudy (anonymous):

That's what I thought, but I was absent this day and didn't know if there was a certain way to do this.

OpenStudy (serenityjustice74):

Oh you have to simplify it as well. Do you know how to break down square roots into simple forms?

OpenStudy (anonymous):

Yeah, don't you have to find factors of them or something?

OpenStudy (serenityjustice74):

Hmm yes, kinda. You get the factors of the number and make sure one of them can be square rooted so you can keep the other inside the square root. @Mehek14 teach please. i feel depressed.

OpenStudy (anonymous):

@ParthKohli

Mehek (mehek14):

we have to simplify √98 and √18 list their factors

OpenStudy (anonymous):

2 * 49 and 2*9

OpenStudy (anonymous):

is that all?

Mehek (mehek14):

so √98 = 7√2 and √18 = 3√2 now we have: √7 + √3 + 7√2- 3√2 we can combine 7√2 - 3√2 just subtract 7 - 3

OpenStudy (anonymous):

alright, and that's it?

Mehek (mehek14):

well tell me what u get :)

OpenStudy (anonymous):

/ = radical /4 + 4/2

Mehek (mehek14):

o-0 what

OpenStudy (anonymous):

idk what im doing

Mehek (mehek14):

all you have to do is combine 7√2 - 3√2 and to do that, subtract 7 - 3 and keep √2 as is

OpenStudy (anonymous):

oh ok so √7 + √3 - 4√2?

Mehek (mehek14):

positive 4 √7 + √3 + 4√2

OpenStudy (anonymous):

yeah, i see why now thanks

Mehek (mehek14):

yw :)

OpenStudy (anonymous):

and what about the other one?

Mehek (mehek14):

3√5 - 3√11 + 2√121 -3√90 we can simplify 2√121 and 3√90 list their factors

OpenStudy (anonymous):

18 * 5?

OpenStudy (anonymous):

wait no

Mehek (mehek14):

no because 18 can be simplified more

OpenStudy (anonymous):

yeah, so 9 and 10

OpenStudy (anonymous):

and 121 is 1 * 121?

Mehek (mehek14):

right so -3√90 = -3*3√10 -3 * 3 = -9 -9√10 and no 121 is a perfect square

OpenStudy (anonymous):

ok, i knew it was a perfect square but i wasnt sure if i had to do anything to it.

Mehek (mehek14):

so 121 just becomes 11 and u had a 2 in front of it so 2 * 11 = ?

OpenStudy (anonymous):

2√11?

Mehek (mehek14):

nope don't need √11 because perfect squares don't need to be radicals it becomes just 11

OpenStudy (anonymous):

oh ok

Mehek (mehek14):

so now multiply 11 * 2

OpenStudy (anonymous):

22

Mehek (mehek14):

so we have 3√5 - 3√11 + 22 -9√10

Mehek (mehek14):

that's all

OpenStudy (anonymous):

alright thanks so much, and sorry for taking so long to understand

Mehek (mehek14):

it's alright :)

OpenStudy (anonymous):

wait, so is that in rational exponent form?

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

you didn't have any exponents tho o-0

OpenStudy (anonymous):

that's what my problem says, my teacher must of made a typo or something

OpenStudy (anonymous):

thanks again

Mehek (mehek14):

yw :)